# Higher Order Dualities between Prime Ideals

### v1

@@ 3 The Prime Ideal Counting Functions and their Estimates
$\alpha>\frac{\log T}{\log Y}=\frac{\log e^{(\log X)^{1-\frac{\delta}{2}}}}{\log e^{(\log X)^{1-\delta}}}=(\log X)^{\frac{\delta}{2}}$ (3.23)
Thus, a standard use of the Stieltjes integral gives us from (3.22) that
$\sum_{\begin{subarray}{c}T<N(\mathfrak{m})\leq\frac{X}{N(\mathfrak{p})^{2}}\\ M_{1}(\mathfrak{m})\leq N(\mathfrak{p})\end{subarray}}\frac{1}{N(\mathfrak{m})}\ll e^{-c_{3}(\log X)^{\frac{\delta}{2}}}$ (3.24)
Therefore, combining (3.17), (3.20), (3.21) and (3.24), we finally have that
$\Psi_{K,2}(X,\mathfrak{p})\ll\frac{X\log N(\mathfrak{p})}{N(\mathfrak{p})\log X}+\frac{X}{N(\mathfrak{p})\log N(\mathfrak{p})}e^{-c_{3}(\log X)^{\frac{\delta}{2}}}$ (3.25)
@@
$\ll\frac{X\log Y}{\log X}$ (3.26)
The last estimate uses a well known number field version of the Merten’s Theorem [Le23] in the first sum, while the second sum is convergent and is much smaller in comparison to the first sum. Thus, we prove our lemma. ∎
@@
$\sum_{\begin{subarray}{c}\mathfrak{p}\subset\mathcal{O}_{K}\\ N(\mathfrak{p})\leq Y\end{subarray}}\Psi_{K,2}(X,\mathfrak{p})\ll\frac{X\log Y}{\log X}$ (3.27)
@@
Therefore, Lemma 3.2 gives us the useful estimate (3.27) that will come in handy to prove the important results in the next section. Before that, we will need another important lemma that gives us a treatment of the number of ideals in $\mathcal{O}_{K}$ contained in at least two prime ideals with the second largest norm, i.e. an estimate on the number of ideals satisfying $Q^{2}(I)\geq 2$. It will be more evident as to why we need this lemma, once we use in the next section (see equation (4.21) in the proof of Theorem 4.1).
@@
$\sum_{\begin{subarray}{c}I\subset\mathcal{O}_{K}\\ 2\leq N(I)\leq X\\ Q^{2}(I)\geq 2\end{subarray}}(Q^{2}(I)-1)\ll\frac{X\log Y}{\log X}+\frac{X\log\log X}{e^{c_{1}\sqrt{\log Y}}}$
@@
$\sum_{\begin{subarray}{c}I\subset\mathcal{O}_{K}\\ 2\leq N(I)\leq X\\ Q^{2}(I)\geq 2\end{subarray}}(Q^{2}(I)-1)\leq\sum_{N(\mathfrak{p})\leq\sqrt{X}}\left|S_{2}\left(\frac{X}{N(\mathfrak{p})},\mathfrak{p}\right)\right|+O\left(\frac{X}{e^{c_{3}\sqrt{\log X\log\log X}}}\right)$ (3.28)
@@
$=\sum_{\begin{subarray}{c}\mathfrak{q}\in\mathcal{O}_{K}\\ \mathfrak{q}\text{ is a prime}\\ N(\mathfrak{p})<N(\mathfrak{q})\leq\frac{X}{N(\mathfrak{p})^{2}}\end{subarray}}\Psi_{K}\left(\frac{X}{N(\mathfrak{p})^{2}N(\mathfrak{q})},N(\mathfrak{p})\right)$ (3.29)
@@
$\sum_{N(\mathfrak{p})\leq Y}\left|S_{2}\left(\frac{X}{N(\mathfrak{p})},\mathfrak{p}\right)\right|\leq\sum_{N(\mathfrak{p})\leq Y}\Psi_{K,2}\left(\frac{X}{N(\mathfrak{p})},\mathfrak{p}\right)\ll\frac{X\log Y}{\log X}$ (3.30)
@@
$\sum_{Y<N(\mathfrak{p})\leq\sqrt{X}}\left|S_{2}\left(\frac{X}{N(\mathfrak{p})},\mathfrak{p}\right)\right|=\sum_{Y<N(\mathfrak{p})\leq\sqrt{X}}\;\sum_{N(\mathfrak{p})<N(\mathfrak{q})\leq\frac{X}{N(\mathfrak{p})^{2}}}\Psi_{K}\left(\frac{X}{N(\mathfrak{p})^{2}N(\mathfrak{q})},N(\mathfrak{p})\right)$ (3.31)
@@
$Y<N(\mathfrak{p}),\quad N(\mathfrak{p})<N(\mathfrak{q}),\quad N(\mathfrak{p})\leq\frac{X}{N(\mathfrak{p})N(\mathfrak{q})},\quad N(\mathfrak{p})\leq\sqrt{X}$
Therefore, with the change in the order of summations and the split, (3.31) results into
@@
$\hskip 56.9055pt+\sum_{\sqrt{X}<N(\mathfrak{q})\leq\frac{X}{Y^{2}}}\;\sum_{\begin{subarray}{c}Y<N(\mathfrak{p})\leq\sqrt{\frac{X}{N(\mathfrak{q})}}\end{subarray}}\Psi_{K}\left(\frac{X}{N(\mathfrak{p})^{2}N(\mathfrak{q})},N(\mathfrak{p})\right)$ (3.32)
We will now use a trick that will not only prove to very useful here in this proof, but also in the proof of Theorem 4.1. This step enables us to use the strong form of the Prime Ideal Theorem (Theorem 1.3) to obtain our estimates888later we will observe that this trick also enables us to involve the strong form of CDT (see Theorem 4.1). Here, we replace the inner sums of both the double sums in the RHS of (3.32) using the following integrals:
$\sum_{\begin{subarray}{c}Y<N(\mathfrak{p})<N(\mathfrak{q})\end{subarray}}\Psi_{K}\left(\frac{X}{N(\mathfrak{p})^{2}N(\mathfrak{q})},N(\mathfrak{p})\right)\quad$ $\text{by}\quad\int_{Y}^{N(\mathfrak{q})}\Psi_{K}\left(\frac{X}{t^{2}N(\mathfrak{q})},t\right)\frac{dt}{\log t}$
$\sum_{\begin{subarray}{c}Y<N(\mathfrak{p})<\sqrt{\frac{X}{N(\mathfrak{q})}}\end{subarray}}\Psi_{K}\left(\frac{X}{N(\mathfrak{p})^{2}N(\mathfrak{q})},N(\mathfrak{p})\right)\quad$ $\text{by}\quad\int_{Y}^{\sqrt{\frac{X}{N(\mathfrak{q})}}}\Psi_{K}\left(\frac{X}{t^{2}N(\mathfrak{q})},t\right)\frac{dt}{\log t}$
@@
$E$ $:=\sum_{\begin{subarray}{c}Y<N(\mathfrak{q})\leq\sqrt{X}\end{subarray}}\left(\sum_{\begin{subarray}{c}Y<N(\mathfrak{p})<N(\mathfrak{q})\end{subarray}}\Psi_{K}\left(\frac{X}{N(\mathfrak{p})^{2}N(\mathfrak{q})},N(\mathfrak{p})\right)-\int_{Y}^{N(\mathfrak{q})}\Psi_{K}\left(\frac{X}{t^{2}N(\mathfrak{q})},t\right)\frac{dt}{\log t}\right)$
$\tilde{E}$ $:=\sum_{\sqrt{X}<N(\mathfrak{q})\leq\frac{X}{Y^{2}}}\left(\sum_{\begin{subarray}{c}Y<N(\mathfrak{p})<\sqrt{\frac{X}{N(\mathfrak{q})}}\end{subarray}}\Psi_{K}\left(\frac{X}{N(\mathfrak{p})^{2}N(\mathfrak{q})},N(\mathfrak{p})\right)-\int_{Y}^{\sqrt{\frac{X}{N(\mathfrak{q})}}}\Psi_{K}\left(\frac{X}{t^{2}N(\mathfrak{q})},t\right)\frac{dt}{\log t}\right)$
@@
$=\sum_{Y<N(\mathfrak{q})\leq\sqrt{X}}\;\sum_{N(\mathfrak{m})\leq\frac{X}{Y^{2}N(\mathfrak{q})}}\left|\sum_{\begin{subarray}{c}\max(Y,M_{1}(\mathfrak{m}))\leq N(\mathfrak{p})\leq\min\left(N(\mathfrak{q}),\sqrt{\frac{X}{N(\mathfrak{mq})}}\right)\end{subarray}}1-\frac{|C|}{|G|}\int_{\max(Y,M_{1}(\mathfrak{m}))}^{\min\left(N(\mathfrak{q}),\sqrt{\frac{X}{N(\mathfrak{mq})}}\right)}\frac{dt}{\log t}\right|$ (3.33)
We observe that the difference inside the the absolute value in (3.33) is almost999we say “almost” due to the truncated intervals of $N(\mathfrak{p})$ exactly the LHS of the strong form of the Prime Ideal Theorem (Theorem 1.3). Note that the error term in Theorem 1.3, i.e. $\frac{X}{e^{c_{1}\sqrt{\log X}}}$ is an increasing function in $X$. Therefore, continuing from (3.33), we can write using Theorem 1.3 that
@@
$\ll\frac{X}{e^{c_{1}\sqrt{\log Y}}}\sum_{Y<N(\mathfrak{q})\leq\sqrt{X}}\frac{1}{N(\mathfrak{q})}\ll\frac{X\log\log X}{e^{c_{1}\sqrt{\log Y}}}$ (3.34)
We can treat the error $\tilde{E}$ in an exact similar manner. We write
@@
$\ll\frac{X}{e^{c_{1}\sqrt{\log Y}}}\sum_{Y<N(\mathfrak{q})\leq\frac{X}{Y^{2}}}\frac{1}{N(\mathfrak{q})}\ll\frac{X\log\log X}{e^{c_{1}\sqrt{\log Y}}}$ (3.35)
Therefore, combining (3.32), (3.34) and (3.35), we get that
@@
$+\sum_{\sqrt{X}<N(\mathfrak{q})\leq\frac{X}{Y^{2}}}\int_{Y}^{\sqrt{\frac{X}{N(\mathfrak{q})}}}\Psi_{K}\left(\frac{X}{t^{2}N(\mathfrak{q})},t\right)\frac{dt}{\log t}+O\left(\frac{X\log\log X}{e^{c_{1}\sqrt{\log Y}}}\right)$ (3.36)
@@
$\beta=u-2;\quad t=e^{\frac{\log\frac{X}{N(\mathfrak{q})}}{u}};\quad du=-\log\frac{X}{N(\mathfrak{q})}\frac{dt}{t(\log t)^{2}}$
@@
$\int_{Y}^{N(\mathfrak{q})}\Psi_{K}\left(\frac{X}{t^{2}N(\mathfrak{q})},t\right)\frac{dt}{\log t}\ll\frac{X}{N(\mathfrak{q})}\int_{2}^{\frac{\log\frac{X}{N(\mathfrak{q})}}{\log Y}}e^{-\frac{\log\frac{X}{N(\mathfrak{q})}}{u}-(u-2)\log(u-2)}\frac{du}{u}\ll\frac{X}{N(\mathfrak{q})Y}$ (3.38)
Hence, the pseudo-integro-sum then can be estimated as
$\sum_{Y<N(\mathfrak{q})\leq\sqrt{X}}\int_{Y}^{N(\mathfrak{q})}\Psi_{K}\left(\frac{X}{t^{2}N(\mathfrak{q})},t\right)\frac{dt}{\log t}\ll\sum_{Y<N(\mathfrak{q})\leq\sqrt{X}}\frac{X}{N(\mathfrak{q})Y}\ll\frac{X\log\log X}{Y}$ (3.39)
@@
$\int_{Y}^{\sqrt{\frac{X}{N(\mathfrak{q})}}}\Psi_{K}\left(\frac{X}{t^{2}N(\mathfrak{q})},t\right)\frac{dt}{\log t}\ll\int_{Y}^{\sqrt{\frac{X}{N(\mathfrak{q})}}}\frac{X}{t^{2}N(\mathfrak{q})}e^{-\beta\log\frac{\beta}{e}}\frac{dt}{\log t}$ (3.40)
@@
$\int_{Y}^{\sqrt{\frac{X}{N(\mathfrak{q})}}}\Psi_{K}\left(\frac{X}{t^{2}N(\mathfrak{q})},t\right)\frac{dt}{\log t}\ll\frac{X}{N(\mathfrak{q})}\int_{2}^{\frac{\log\frac{X}{N(\mathfrak{q})}}{\log Y}}e^{-\frac{\log\frac{X}{N(\mathfrak{q})}}{u}-(u-2)\log(u-2)}\frac{du}{u}\ll\frac{X}{N(\mathfrak{q})Y}$ (3.41)
Therefore, again the second pseudo-integro-sum in (3.36) can be estimated as
$\sum_{\sqrt{X}<N(\mathfrak{q})\leq\frac{X}{Y^{2}}}\int_{Y}^{\sqrt{\frac{X}{N(\mathfrak{q})}}}\Psi_{K}\left(\frac{X}{t^{2}N(\mathfrak{q})},t\right)\frac{dt}{\log t}\ll\frac{X\log\log X}{Y}$ (3.42)
Therefore, finally combining (3.36), (3.39) and (3.42), we get
$\sum_{\begin{subarray}{c}Y<N(\mathfrak{p})\leq\sqrt{X}\end{subarray}}\left|S_{2}\left(\frac{X}{N(\mathfrak{p})},\mathfrak{p}\right)\right|\ll\frac{X\log\log X}{Y}+\frac{X\log\log X}{e^{c_{1}\sqrt{\log Y}}}\ll\frac{X\log\log X}{e^{c_{1}\sqrt{\log Y}}}$ (3.43)
Thus, combining (3.28), (3.30) and (3.43) together yield our required result. ∎
@@
Note that we have deliberately kept two different error terms dependent on the choice of $Y$. We will make a suitable choice of $Y$ in the proof of Theorem 4.1 (see below) where we will use Lemma 3.4.

### v2

@@ 3 The Prime Ideal Counting Functions and their Estimates
$\alpha>\frac{\log T}{\log Y}=\frac{\log e^{(\log X)^{1-\frac{\delta}{2}}}}{\log e^{(\log X)^{1-\delta}}}=(\log X)^{\frac{\delta}{2}}.$ (3.23)
Thus, from (3.22) and (3.23), we get that
$\sum_{\begin{subarray}{c}T<N(\mathfrak{m})\leq\frac{X}{N(\mathfrak{p})^{2}}\\ M_{1}(\mathfrak{m})\leq N(\mathfrak{p})\end{subarray}}\frac{1}{N(\mathfrak{m})}\ll e^{-c_{3}(\log X)^{\frac{\delta}{2}}}.$ (3.24)
Therefore, combining (3.17), (3.20), (3.21), and (3.24), we finally have that
$\Psi_{K,2}(X,\mathfrak{p})\ll\frac{X\log N(\mathfrak{p})}{N(\mathfrak{p})\log X}+\frac{X}{N(\mathfrak{p})\log N(\mathfrak{p})}e^{-c_{3}(\log X)^{\frac{\delta}{2}}}.$ (3.25)
@@
$\ll\frac{X\log Y}{\log X}.$ (3.26)
The last estimate uses a well-known number field version of Merten’s Theorem [Le23] in the first sum, while the second sum is convergent and is much smaller in comparison to the first sum. Thus, we prove our lemma. ∎
@@
$\sum_{\begin{subarray}{c}\mathfrak{p}\subset\mathcal{O}_{K}\\ N(\mathfrak{p})\leq Y\end{subarray}}\Psi_{K,2}(X,\mathfrak{p})\ll\frac{X\log Y}{\log X}.$ (3.27)
@@
Therefore, Lemma 3.2 gives us the useful estimate (3.27) that will come in handy to prove the important results in the next section. Before that, we will need another important lemma that gives us a treatment of the number of ideals in $\mathcal{O}_{K}$ contained in at least two prime ideals with the second largest norm, i.e., an estimate on the number of ideals satisfying $Q^{2}(I)\geq 2$. It will be more evident as to why we need this lemma, once we use it in the next section (see equation (4.21) in the proof of Theorem 4.1).
@@
$\sum_{\begin{subarray}{c}I\subset\mathcal{O}_{K}\\ 2\leq N(I)\leq X\\ Q^{2}(I)\geq 2\end{subarray}}(Q^{2}(I)-1)\ll\frac{X\log Y}{\log X}+\frac{X\log\log X}{e^{c_{1}\sqrt{\log Y}}}.$
@@
$\sum_{\begin{subarray}{c}I\subset\mathcal{O}_{K}\\ 2\leq N(I)\leq X\\ Q^{2}(I)\geq 2\end{subarray}}(Q^{2}(I)-1)\leq\sum_{N(\mathfrak{p})\leq\sqrt{X}}\left|S_{2}\left(\frac{X}{N(\mathfrak{p})},\mathfrak{p}\right)\right|+O\left(\frac{X}{e^{c_{3}\sqrt{\log X\log\log X}}}\right).$ (3.28)
@@
$=\sum_{\begin{subarray}{c}\mathfrak{q}\in\mathcal{O}_{K}\\ \mathfrak{q}\text{ is a prime}\\ N(\mathfrak{p})<N(\mathfrak{q})\leq\frac{X}{N(\mathfrak{p})^{2}}\end{subarray}}\Psi_{K}\left(\frac{X}{N(\mathfrak{p})^{2}N(\mathfrak{q})},N(\mathfrak{p})\right).$ (3.29)
@@
$\sum_{N(\mathfrak{p})\leq Y}\left|S_{2}\left(\frac{X}{N(\mathfrak{p})},\mathfrak{p}\right)\right|\leq\sum_{N(\mathfrak{p})\leq Y}\Psi_{K,2}\left(\frac{X}{N(\mathfrak{p})},\mathfrak{p}\right)\ll\frac{X\log Y}{\log X}.$ (3.30)
@@
$\sum_{Y<N(\mathfrak{p})\leq\sqrt{X}}\left|S_{2}\left(\frac{X}{N(\mathfrak{p})},\mathfrak{p}\right)\right|=\sum_{Y<N(\mathfrak{p})\leq\sqrt{X}}\;\sum_{N(\mathfrak{p})<N(\mathfrak{q})\leq\frac{X}{N(\mathfrak{p})^{2}}}\Psi_{K}\left(\frac{X}{N(\mathfrak{p})^{2}N(\mathfrak{q})},N(\mathfrak{p})\right).$ (3.31)
@@
$Y<N(\mathfrak{p}),\quad N(\mathfrak{p})<N(\mathfrak{q}),\quad N(\mathfrak{p})\leq\frac{X}{N(\mathfrak{p})N(\mathfrak{q})},\quad N(\mathfrak{p})\leq\sqrt{X}.$
Therefore, with the change in the order of summations and the split, (3.31) results in
@@
$\hskip 56.9055pt+\sum_{\sqrt{X}<N(\mathfrak{q})\leq\frac{X}{Y^{2}}}\;\sum_{\begin{subarray}{c}Y<N(\mathfrak{p})\leq\sqrt{\frac{X}{N(\mathfrak{q})}}\end{subarray}}\Psi_{K}\left(\frac{X}{N(\mathfrak{p})^{2}N(\mathfrak{q})},N(\mathfrak{p})\right).$ (3.32)
We will now use a trick that will not only prove to be very useful here in this proof, but also in the proof of Theorem 4.1. This step enables us to use the strong form of the Prime Ideal Theorem (Theorem 1.3) to obtain our estimates888later we will observe that this trick also enables us to involve the strong form of CDT (see Theorem 4.1). Here, we replace the inner sums of both the double sums in the RHS of (3.32) using the following integrals:
$\sum_{\begin{subarray}{c}Y<N(\mathfrak{p})<N(\mathfrak{q})\end{subarray}}\Psi_{K}\left(\frac{X}{N(\mathfrak{p})^{2}N(\mathfrak{q})},N(\mathfrak{p})\right)\quad$ $\text{by}\quad\int_{Y}^{N(\mathfrak{q})}\Psi_{K}\left(\frac{X}{t^{2}N(\mathfrak{q})},t\right)\frac{dt}{\log t},$
$\sum_{\begin{subarray}{c}Y<N(\mathfrak{p})<\sqrt{\frac{X}{N(\mathfrak{q})}}\end{subarray}}\Psi_{K}\left(\frac{X}{N(\mathfrak{p})^{2}N(\mathfrak{q})},N(\mathfrak{p})\right)\quad$ $\text{by}\quad\int_{Y}^{\sqrt{\frac{X}{N(\mathfrak{q})}}}\Psi_{K}\left(\frac{X}{t^{2}N(\mathfrak{q})},t\right)\frac{dt}{\log t}.$
@@
$E$ $:=\sum_{\begin{subarray}{c}Y<N(\mathfrak{q})\leq\sqrt{X}\end{subarray}}\left(\sum_{\begin{subarray}{c}Y<N(\mathfrak{p})<N(\mathfrak{q})\end{subarray}}\Psi_{K}\left(\frac{X}{N(\mathfrak{p})^{2}N(\mathfrak{q})},N(\mathfrak{p})\right)-\int_{Y}^{N(\mathfrak{q})}\Psi_{K}\left(\frac{X}{t^{2}N(\mathfrak{q})},t\right)\frac{dt}{\log t}\right),$
$\tilde{E}$ $:=\sum_{\sqrt{X}<N(\mathfrak{q})\leq\frac{X}{Y^{2}}}\left(\sum_{\begin{subarray}{c}Y<N(\mathfrak{p})<\sqrt{\frac{X}{N(\mathfrak{q})}}\end{subarray}}\Psi_{K}\left(\frac{X}{N(\mathfrak{p})^{2}N(\mathfrak{q})},N(\mathfrak{p})\right)-\int_{Y}^{\sqrt{\frac{X}{N(\mathfrak{q})}}}\Psi_{K}\left(\frac{X}{t^{2}N(\mathfrak{q})},t\right)\frac{dt}{\log t}\right).$
@@
$=\sum_{Y<N(\mathfrak{q})\leq\sqrt{X}}\;\sum_{N(\mathfrak{m})\leq\frac{X}{Y^{2}N(\mathfrak{q})}}\left|\sum_{\begin{subarray}{c}\max(Y,M_{1}(\mathfrak{m}))\leq N(\mathfrak{p})\leq\min\left(N(\mathfrak{q}),\sqrt{\frac{X}{N(\mathfrak{mq})}}\right)\end{subarray}}1-\int_{\max(Y,M_{1}(\mathfrak{m}))}^{\min\left(N(\mathfrak{q}),\sqrt{\frac{X}{N(\mathfrak{mq})}}\right)}\frac{dt}{\log t}\right|.$ (3.33)
We observe that the difference inside the absolute value in (3.33) is almost999we say “almost” due to the truncated intervals of $N(\mathfrak{p})$ exactly the LHS of the strong form of the Prime Ideal Theorem (Theorem 1.3). Note that the error term in Theorem 1.3, i.e. $\frac{X}{e^{c_{1}\sqrt{\log X}}}$ is an increasing function in $X$. Therefore, continuing from (3.33), we can write using Theorem 1.3 that
@@
$\ll\frac{X}{e^{c_{1}\sqrt{\log Y}}}\sum_{Y<N(\mathfrak{q})\leq\sqrt{X}}\frac{1}{N(\mathfrak{q})}\ll\frac{X\log\log X}{e^{c_{1}\sqrt{\log Y}}}.$ (3.34)
We can treat the error $\tilde{E}$ in an exactly similar manner. We write
@@
$\ll\frac{X}{e^{c_{1}\sqrt{\log Y}}}\sum_{Y<N(\mathfrak{q})\leq\frac{X}{Y^{2}}}\frac{1}{N(\mathfrak{q})}\ll\frac{X\log\log X}{e^{c_{1}\sqrt{\log Y}}}.$ (3.35)
Therefore, combining (3.32), (3.34), and (3.35), we get that
@@
$+\sum_{\sqrt{X}<N(\mathfrak{q})\leq\frac{X}{Y^{2}}}\int_{Y}^{\sqrt{\frac{X}{N(\mathfrak{q})}}}\Psi_{K}\left(\frac{X}{t^{2}N(\mathfrak{q})},t\right)\frac{dt}{\log t}+O\left(\frac{X\log\log X}{e^{c_{1}\sqrt{\log Y}}}\right).$ (3.36)
@@
$\beta=u-2;\quad t=e^{\frac{\log\frac{X}{N(\mathfrak{q})}}{u}};\quad du=-\log\frac{X}{N(\mathfrak{q})}\frac{dt}{t(\log t)^{2}}.$
@@
$\int_{Y}^{N(\mathfrak{q})}\Psi_{K}\left(\frac{X}{t^{2}N(\mathfrak{q})},t\right)\frac{dt}{\log t}\ll\frac{X}{N(\mathfrak{q})}\int_{2}^{\frac{\log\frac{X}{N(\mathfrak{q})}}{\log Y}}e^{-\frac{\log\frac{X}{N(\mathfrak{q})}}{u}-(u-2)\log(u-2)}\frac{du}{u}\ll\frac{X}{N(\mathfrak{q})Y}.$ (3.38)
Hence, the pseudo-integro-sum can then be estimated as
$\sum_{Y<N(\mathfrak{q})\leq\sqrt{X}}\int_{Y}^{N(\mathfrak{q})}\Psi_{K}\left(\frac{X}{t^{2}N(\mathfrak{q})},t\right)\frac{dt}{\log t}\ll\sum_{Y<N(\mathfrak{q})\leq\sqrt{X}}\frac{X}{N(\mathfrak{q})Y}\ll\frac{X\log\log X}{Y}.$ (3.39)
@@
$\int_{Y}^{\sqrt{\frac{X}{N(\mathfrak{q})}}}\Psi_{K}\left(\frac{X}{t^{2}N(\mathfrak{q})},t\right)\frac{dt}{\log t}\ll\int_{Y}^{\sqrt{\frac{X}{N(\mathfrak{q})}}}\frac{X}{t^{2}N(\mathfrak{q})}e^{-\beta\log\frac{\beta}{e}}\frac{dt}{\log t}.$ (3.40)
@@
$\int_{Y}^{\sqrt{\frac{X}{N(\mathfrak{q})}}}\Psi_{K}\left(\frac{X}{t^{2}N(\mathfrak{q})},t\right)\frac{dt}{\log t}\ll\frac{X}{N(\mathfrak{q})}\int_{2}^{\frac{\log\frac{X}{N(\mathfrak{q})}}{\log Y}}e^{-\frac{\log\frac{X}{N(\mathfrak{q})}}{u}-(u-2)\log(u-2)}\frac{du}{u}\ll\frac{X}{N(\mathfrak{q})Y}.$ (3.41)
Therefore, again, the second pseudo-integro-sum in (3.36) can be estimated as
$\sum_{\sqrt{X}<N(\mathfrak{q})\leq\frac{X}{Y^{2}}}\int_{Y}^{\sqrt{\frac{X}{N(\mathfrak{q})}}}\Psi_{K}\left(\frac{X}{t^{2}N(\mathfrak{q})},t\right)\frac{dt}{\log t}\ll\frac{X\log\log X}{Y}.$ (3.42)
Therefore, finally combining (3.36), (3.39), and (3.42), we get
$\sum_{\begin{subarray}{c}Y<N(\mathfrak{p})\leq\sqrt{X}\end{subarray}}\left|S_{2}\left(\frac{X}{N(\mathfrak{p})},\mathfrak{p}\right)\right|\ll\frac{X\log\log X}{Y}+\frac{X\log\log X}{e^{c_{1}\sqrt{\log Y}}}\ll\frac{X\log\log X}{e^{c_{1}\sqrt{\log Y}}}.$ (3.43)
Thus, combining (3.28), (3.30), and (3.43) together yields our required result. ∎
@@
Note that we have deliberately kept two different error terms dependent on the choice of $Y$. We will make a suitable choice of $Y$ in the proof of Theorem 4.1 (see below), where we will use Lemma 3.4.